OpenStudy (anonymous):

What are 5 different mixed numbers that when added together equal 10? This is an emergency! WILL MEDAL AND FAN! The fractions can only be 1/4 or/and 1/3.

OpenStudy (nysha):

I'm sorry don't know

OpenStudy (anonymous):

1.2.5.10

OpenStudy (anonymous):

1x10=10 2x5=10 5x2=10 10x1=10

OpenStudy (anonymous):

\[\frac{ 1 }{ 1 } \frac{ 2 }{ 2 }\frac{ 5 }{ 5 }\frac{ 10 }{ 10 }\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

well i dont know what to tell you !!

OpenStudy (anonymous):

@Hero @jigglypuff314